# Ladder-residual: parallelism-aware architecture for accelerating large model inference with communication overlapping

### v2

@@ 3.3.2 Benchmarking
Figure 3: End-to-end inference throughput improvement on Llama-3-405B on a generation task with $1024$ prompt tokens and $512$ completion tokens. Here we use TP size 16 across two nodes each with 8 GPUs, connected with InfiniBand. “Upper Bound” denotes the model without doing communication at all (which results in incorrect computation). Due to the high cost of cross-node communication, Ladder Residual architecture is able to achieve more than $30\%$ improvement across various batch sizes with NVLink enabled and around $50\%$ without NVLink.
@@
First, we measure the best latency achievable using both the Ladder Residual architecture and a traditional transformer baseline, both in terms of end-to-end latency and broken down by inference phase (prefill vs decode). The best latency in our setting is achieved using a batch size of $1$ and a TP degree of $8$ (the maximum TP world size possible on our $8$-GPU node). We present the results of these latency-optimized experiments in Table 2. In this latency-optimized regime, both with and without NVLink, we find that the Ladder Residual architecture outperforms the parallel attn-mlp alternative in both prefill and decode latency.
Second, we measure the throughput at each TP world size, and at a representative selection of batch sizes, for 70B model. We present the results of these experiments in Figure 2. In these throughput-oriented experiments, we again find that the Ladder Residual architecture significantly outperforms the traditional transformer architecture, and that these improvements are robust to the presence or absence of NVLink. We also find that the throughput gains from adopting the Ladder Residual architecture increase as the TP degree increases, reflecting the greater proportion of run time spent in communication relative to compute as we partition the computation across a larger number of devices. Lastly, as shown in Table 1, the amount of improvement decrease as we increase from 8B to 70B when running with NVLink, as the computation scales faster than communication. However, the trend is reversed when running without NVLink, likely due to the much higher cost of communication in that scenario.
Serving a model of larger size can be a challenge since even loading the model can require more than 8 GPUs. Cross-node TP communication is very expensive, however the common practice of using intra-node TP with cross-node Pipeline Parallelism (PP) is dependent on batch size to reduce gpu idle time (for example, with batch size = 1, half of the GPUs will be idle at any given time). With the speedup from Ladder Residual, cross-node TP can be a viable option. We benchmarked 405B under such setting in Figure 3, found that even for nodes with fast InfiniBand interconnect, Ladder Residual architecture can achieve more than $30\%$ throughput improvement across various batch sizes.
@@ 4.1 Training From Scratch
We train a 1B and 3B Ladder Transformer from scratch and compare its performance with an equally sized standard Transformer model. All our models are trained on 100B tokens of FineWeb-edu dataset (HuggingFaceFW, 2024) using the StarCoder tokenizer (Li et al., 2023b). We also compare our model with the Parallel Attention/MLP architecture (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021) which parallelizes the computation of the attention and the MLP module. This effectively reduces the communication cost by 50% for the tensor parallel all-reduce in both the forward and backward computation.
@@ 4.1.1 Experimental details
We use DDP (Distributed Data Parallel) (Li et al., 2020) to train the 1B and HSDP (Hybrid Sharded Data Parallel) (Zhao et al., 2023; Rajbhandari et al., 2020) to train the 3B models. For HSDP, we shard the model within 1 node (equipped with 8x H100 GPUs) and replicate the model outside the node. We use mixed precision training (Micikevicius et al., 2018) in BF16 (Kalamkar et al., 2019) with gradient accumulation and gradient all-reduce/reduce-scatter in FP32 for training stability. We train all our models with 2048 context length with a batch size of 4M tokens in a batch. The models are trained with cosine scheduler with a warmup of 8B tokens to a peak learning rate of $3\times 10^{-4}$. The learning rate is then decayed over 92B tokens to $3\times 10^{-5}$.
We use EleutherAI’s LM eval harness (Gao et al., 2024) to evaluate all our models on ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017) and Winogrande (Trinh & Le, 2018). We also evaluate Wikitext (Merity et al., 2017) perplexity for all our models.
@@ 4.1.2 Results
The full results can be found at Table 3. We find that at the 1B model scale, the Ladder Transformer achieves obtains similar performance compared with the Standard Transformer while beating the Parallel Transformer. At the 3B parameter scale, we find that the Standard Transformer baseline model is better than the Ladder Transformer model with 3.2% lower perplexity and 1.2 points of absolute difference in accuracy. The Parallel Transformer has almost the same performance as the Ladder Transformer at the 3B scale.
@@ 4.2 Post-training adaptation
We investigate the feasibility of directly applying Ladder Residual on an existing pre-trained model, and we choose Llama-3.1 8B Instruct as the target. We applied Ladder Residual to the upper half of the Llama to keep the performance since we hypothesize that touching the lower layers can destroy knowledge that is hard to recover without large-scale retraining. We evaluate the adapted models on 8 benchmarks across a range of domains: accuracy on MMLU (5-shots) (Hendrycks et al., 2021) and ARC-Challenge (ARC-C, 25-shots) (Clark et al., 2018), normalized accuracy on OpenBookQA (OBQA) (Mihaylov et al., 2018), HellaSwag (HS, 10-shots) (Zellers et al., 2019), and TruthfulQA (TQ, mc1) (Lin et al., 2022). exact-match accuracy on GSM8K(GSM, 8-shots) (Cobbe et al., 2021), pass@1 on HumanEval+(HE+) (Chen et al., 2021), aggregated accuracy on IFEval (Zhou et al., 2023), and length controlled win rate (Dubois et al., 2024) against gpt4-turbo on AlpacaEval (Li et al., 2023c). The evaluation of HumanEval+ is conducted with EvalPlus (Liu et al., 2023a), AlpacaEval is done with the AlpacaEval2 library, and the rest of the evaluations are conducted with the LM-Evaluation-Harness library (Gao et al., 2024).
@@ 4.2.1 Experimental details
We convert a state-of-the-art open-source model, Llama-3.1-8B-Instruct into a hybrid Ladder Residual structure, by applying Ladder Residual to the upper half of the model (layers 16-32 for LLaMA-3.1-8B-Instruct). We call this variant Hybrid-Ladder-8B-16L in Table 4. We also experiment with more aggressive adaptation where we applied Ladder Residual to the layers 12-32 and we call this experiment Hybrid-Ladder-8B-20L. We conduct supervised fine-tuning (SFT) for the resulting model on the 7M subset and the Gen subset of the Infinity-Instruct dataset***https://huggingface.co/datasets/BAAI/Infinity-Instruct, which contains 3B tokens. We train for 2 epochs with AdamW optimizer with a batch size of 32. We use 5e-6 learning rate with 200 steps of linear warmup, followed by cosine annealing to the end. We use Axolotl ***https://github.com/axolotl-ai-cloud/axolotl for our SFT experiments.
@@
As shown in Table 4, after adaptation, there is a huge performance drop mainly on generative tasks as the computation flow is messed up. But after light retraining, the hybrid Ladder Llama is able to reach the same level of performance with the original Llama. By applying Ladder Residual on the last 16 layers (half of the 32 layers), we are able to obtain 14.5% end-to-end wall clock speed up at the inference time with TP world size of 8 and batch size of 1. Our results demonstrate the potential of Ladder Residual being a drop-in adaptation technique to make the model faster without sacrificing performance. We additionally experiment with applying Ladder Residual to the last 20 layers of Llama and found that it leads to a slight drop in performance. There is a chance that with longer adaptation, or smarter adaptation techniques like distillation or iterative training, we can obtain a Ladder-Llama with more layers adapted. We leave the further exploration to future work.
5 Eliminating communication with Desynced Residual
Figure 5: Illustration of Desynced Residual-2x on the left and Desynced Residual-4x on the right. Each GPU processes its local hidden states with reduced communication. In Desynced Residual-2x, the attention output is not synchronized and each GPU will have its own attention output to be processed by the next MLP module whereas in Desynced Residual-4x the activations are synchonized after the MLP of every alternate layer. Note that the different colors for the residual stream indicate the desynchronization of the residual stream here which is re-synchronized after we call an AllReduce operation.
We find that in settings where there is no NVLink connection, the latency for AllReduce dominates and its not possible to hide the latency of the AllReduce completely with the computation of another block in the Ladder Residual method. We thus propose an alternative method, Desynced Residual, that entirely drops the AllReduce communication and lets each device process its own activations independently. We call this architecture Desynced Residual because dropping the communication leads to desynchronization of the residual stream in the model which is re-synchronized at the next AllReduce operation. We experiment with 2 different Desync configurations: Desync Residual-2x and Desync Residual-4x. Desync Residual-$n$x means that we only retain the last AllReduce operation in a group of $n$ sequential AllReduce operations i.e ($n-1$) AllReduce operations are dropped from the model. While it is possible to use an arbitrary communication pattern, we find that dropping AllReduce for Attention yields a model with lower Wikitext perplexity than dropping AllReduce for MLP while offering a good tradeoff between inference speed and model accuracy. We compare the latency improvement of Desync Residual-nx and Ladder Residual in Table 6. While Desync Residual-2x performs slightly worse than Ladder Residual, Desynced Residual-4x is able to achieve a larger gain, especially under the no NVLink setup ($39\%$ for Desynced Residual-4x compared to $23\%$ for Ladder Residual). The promising speedup shows the potential of Desync Residual as an alternative architecture to be further explored especially in cases where communication can be very expensive. Figure 5 shows an illustration of the Desynced Residual-2x and Desynced Residual-4x model running on 2 GPUs.
5.1 Experiments and Results
We run both 1B and 3B pretraining from scratch experiments for Desync Residual-2x and Desync Residual-4x models in the same setting as described in subsection 4.1. We find that the Desync Residual-4x model performs better than Desync Residual-2x model on 1B scale on Wikitext perplexity while being slightly worse on accuracy. While on the 3B scale, the Desync Residual-4x model is better on both perplexity and average accuracy. This is quite surprising and demonstrates that it is possible to train models using the Desync Residual in a way to significantly reduce communication. For instance, the Desync Residual-4x drops 75% communications with almost no drop in model performance.
In our benchmarking results, we find that at a large batch sizes (64 in our experiments) with a TP world size of 8, it is possible to achieve around 30% improvement in inference throughput over NVLink and around 40% improvement without NVLink using the Desync Residual-4x architecture. We also observe a significant improvement in the first token latency (prefill latency) using the Desync Residual-4x model over Ladder Residual and Desync Residual-2x. The first token latency is especially important in scenarios where the LLM might be used as a classifier for example or in cases where multiple calls to the LLM need to be issued (agentic usecases for example). By manipulating where to communicate, Desync Residual allows designing networks that can drop arbitrary amount of communication, which allows flexible architectural design giving different needs. For setup with very slow communication (for example wireless communication), Desync Residual enables sacrificing a small amount of performance in exchange for speed.
Table 5: Performance of Desync Residual compared to the Standard Transformer model trained on FineWeb-edu for 100B tokens.
Desync Residual-2x-1B 32.51 69.36 40.53 72.03 86.00 56.04 59.41 18.70
Desync Residual-4x-1B 32.17 68.60 41.25 71.60 87.40 54.70 59.29 18.58
Desync Residual-2x-3B 37.97 73.57 46.56 74.43 90.30 58.80 63.61 14.67
Desync Residual-4x-3B 38.57 73.15 46.49 73.72 91.90 58.64 63.75 14.60
Table 6: Detailed breakdown for prefill latency, decode latency and token/sec improvement (%) for different models. The speedup (%) is calculated by using the latency of the optimized model divided by the original model. All experiments are based on batch size 64, TP degree 8.
Model Prefill Improvement (%) Decode Improvement (%) Token/sec Improvement (%)
NVLINK-UpperBound-Llama-8B 25.43 27.22 34.31
NVLINK-Ladder-Llama-8B 9.93 18.70 14.43
NVLINK-Desync-Residual-2x-Llama-8B 8.40 12.03 10.51
NVLINK-Desync-Residual-4x-Llama-8B 17.07 16.53 20.36
NO-NVLINK-UpperBound-Llama-8B 31.60 51.95 64.97
NO-NVLINK-Ladder-Llama-8B 15.07 26.26 23.98
NO-NVLINK-Desync-Residual-2x-Llama-8B 13.36 24.36 21.59
NO-NVLINK-Desync-Residual-4x-Llama-8B 22.11 37.73 39.01
@@ 6 Related Work
Overlapping communication has been a widely explored area in prior works in order to achieve higher performance for distributed training. For Tensor Parallelism, prior works (Jangda et al., 2022; Wang et al., 2022; NVIDIA, 2023) decompose the communication into more fine-grained operations in order to find computations with no dependency to overlap. Our work doesn’t rely on such decompositions and therefore doesn’t require Sequence Parallelism to handle the partitioned activations before all-gather. In FSDP (FairScale authors, 2021), the all-gather communication is usually prefetched to be overlapped with the communication. Pipeline Parallelism (NVIDIA, 2023; Li et al., 2023a; Lamy-Poirier, 2023) on the other hand, chunks the data into mini-batches which creates more opportunity for overlapping. Compared to these other parallelism approaches, TP has the advantage to be independent from the batch size or sequence length, and is able to partition the computation as much as possible given enough GPUs in theory.
@@
Prior works have explored various alternative designs for Transformer, for example parallel attention and mlp (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021), linear attention (Katharopoulos et al., 2020), Grouped Query Attention (Ainslie et al., 2023), Cross-Layer Attention (Brandon et al., 2024) to improve the training and inference efficiency. Some of these variants are more widely adopted than others, due to the degree of impact they have on performance and efficiency. Past works have also considered adapting an existing checkpoint to these efficient variants.  Ainslie et al. (2023) extracted grouped-query attention from a multi-head attention model, and  Wang et al. (2024) converted a Llama model to a Mamba model by retraining on 50B tokens to close the performance gap. Comparatively, our adaptation is much lighter (3B tokens), showing that the representation shift introduced by Ladder Residual is easier to recover.  Wang et al. (2024) considered converting a Llama model to a Mamba model and used distillation to retrain the converted model. Such training paradigms that specifically tune the model to align with the original model could further improve the Ladder Residual based models.
@@ 7 Conclusion
We introduce Ladder Residual and Desync Residual, architectural modifications that allow overlapping communication with computation or dropping it entirely for model parallelism. We show that when running Tensor Parallelism, Ladder Residual can achieve great speed up across various model sizes, batch sizes, and number of GPUs. When applying Ladder Residual to Llama-3.1 8B Instruct, we only need lightweight retraining to reach the same level of performance with the original model, showing its potential to be a plug-in for any pretrained Transformer and we are able to obtain 23% end-to-end wall clock speed up at the inference time with TP world size of 8 and batch size of 1. We also trained a 1B and 3B Ladder Transformer from scratch, and found that they are comparable to the standard Transformer of the same size while achieving over 55% speedup. On the other hand, Desync Residual provides flexible removal of communication when designing an architecture to train from scratch, which can be helpful in the case where the communication is too costly to overlap. Given that such a simple architectural change can obviate the need for expensive interconnects while maintaining model quality, we hope that our methods will inspire even closer co-design between model architecture and inference systems.

### v4

@@ 3.3.2 Benchmarking
Figure 3: End-to-end inference throughput improvement on Llama-3-405B on a generation task with $1024$ prompt tokens and $512$ completion tokens. Here we use TP size 16 across two nodes each with 8 GPUs, connected with InfiniBand. Due to the high cost of cross-node communication, Ladder Residual architecture is able to achieve more than $30\%$ improvement across various batch sizes with P2P enabled and around $50\%$ with P2P communication disabled.
@@
First, we measure the best latency achievable (batch size 1, TP degree 8) using both the Ladder Residual architecture and a standard transformer baseline. We report end-to-end latency broken down by inference phase (prefill vs decode) in Table 2. In this latency-optimized regime, both with and without NVLink, Ladder Residual outperforms the parallel attn-mlp alternative in both prefill and decode latency.
Second, we measure the throughput across different TP world size and batch sizes for 70B model in Figure 2. In these throughput-oriented experiments, we again find that the Ladder Residual architecture significantly outperforms the standard transformer architecture, and that these improvements are larger when communication is slower. The throughput gains from adopting the Ladder Residual architecture increase as the TP degree increases, reflecting the greater proportion of run time spent in communication relative to compute as we partition the computation across a larger number of devices. Lastly, as shown in Table 1, the amount of improvement decrease as we increase from 8B to 70B when running with P2P communication, as the computation scales faster than communication. However, the trend is reversed when running without P2P communication, due to the much higher cost of communication in that scenario.
Finally, we consider serving model that is too large to be loaded on one node. Cross-node TP communication is very expensive, however the common practice of using intra-node TP with cross-node Pipeline Parallelism (PP) is dependent on batch size to reduce gpu idle time (for example, with batch size = 1, half of the GPUs will be idle at any given time). With the speedup from Ladder Residual, cross-node TP can be a viable option. We benchmarked 405B under such setting in Figure 3, found that even for nodes with fast InfiniBand interconnect, Ladder Residual can achieve more than $30\%$ throughput improvement across batch sizes.
@@ 4.1 Training From Scratch
We train a 1B and 3B Ladder Transformer from scratch and compare its performance with an equally sized standard Transformer model. All our models are trained on 100B tokens of FineWeb-edu dataset (HuggingFaceFW, 2024) using the StarCoder tokenizer (Li et al., 2023b). We also compare our model with the Parallel Attention/MLP architecture (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021) which parallelizes the computation of the attention and the MLP module. This effectively reduces the communication cost by 50% for the tensor parallel AllReduce in both the forward and backward computation.
@@ 4.1.1 Experimental details
We use DDP (Distributed Data Parallel) (Li et al., 2020) to train the 1B and HSDP (Hybrid Sharded Data Parallel) (Zhao et al., 2023; Rajbhandari et al., 2020) to train the 3B models. For HSDP, we shard the model within 1 node (equipped with 8x H100 GPUs) and replicate the model outside the node. We use mixed precision training (Micikevicius et al., 2018) in BF16 (Kalamkar et al., 2019) with gradient accumulation and gradient AllReduce/ReduceScatter in FP32 for training stability. We train all our models with 2048 context length with a batch size of 4M tokens in a batch. The models are trained with cosine scheduler with a warmup of 8B tokens to a peak learning rate of $3\times 10^{-4}$. The learning rate is then decayed over 92B tokens to $3\times 10^{-5}$.
We use EleutherAI’s LM eval harness (Gao et al., 2024) to evaluate models on ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017) and Winogrande (Trinh & Le, 2018). We also evaluate perplexity on Wikitext (Merity et al., 2017).
@@ 4.1.2 Results
The full results can be found at Table 3. We find that at the 1B model scale, Ladder Transformer achieves performance similar to Standard Transformer while beating the Parallel Transformer. At 3B parameter scale, we find that Standard Transformer is better than Ladder Transformer model with 3.2% lower perplexity and 1.2 points of absolute difference in accuracy. The Parallel Transformer has almost the same performance as Ladder Transformer at the 3B scale.
@@ 4.2 Post-training adaptation
We investigate the feasibility of directly applying Ladder Residual on an existing pre-trained model. We applied Ladder Residual to the upper half of the Llama-3.1 8B Instruct to keep the performance since we found that touching the lower layers can destroy knowledge that is hard to recover without large-scale retraining. We evaluate the adapted models on 8 benchmarks across a range of domains: accuracy on MMLU (5-shots) (Hendrycks et al., 2021) and ARC-Challenge (ARC-C, 25-shots) (Clark et al., 2018), normalized accuracy on OpenBookQA (OBQA) (Mihaylov et al., 2018), HellaSwag (HS, 10-shots) (Zellers et al., 2019), and TruthfulQA (TQ, mc1) (Lin et al., 2022). exact-match accuracy on GSM8K(GSM, 8-shots) (Cobbe et al., 2021), pass@1 on HumanEval+(HE+) (Chen et al., 2021), aggregated accuracy on IFEval (Zhou et al., 2023), and length controlled win rate (Dubois et al., 2024) against gpt4-turbo on AlpacaEval (Li et al., 2023c). The evaluation of HumanEval+ is conducted with EvalPlus (Liu et al., 2023a), AlpacaEval is done with the AlpacaEval2 library, and the rest of the evaluations are conducted with the LM-Evaluation-Harness library (Gao et al., 2024).
@@ 4.2.1 Experimental details
We convert a state-of-the-art open-source model, Llama-3.1-8B-Instruct into a hybrid Ladder Residual structure, by applying Ladder Residual to the upper half of the model (layers 16-32 for LLaMA-3.1-8B-Instruct). We call this variant Hybrid-Ladder-8B-16L in Table 4. We also experiment with more aggressive adaptation where we applied Ladder Residual to the layers 12-32 and we call this experiment Hybrid-Ladder-8B-20L. We conduct supervised fine-tuning (SFT) for the resulting model on the 7M subset and the Gen subset of the Infinity-Instruct dataset [^3] https://huggingface.co/datasets/BAAI/Infinity-Instruct, which contains 3B tokens. We train for 2 epochs with AdamW optimizer with a batch size of 32. We use $5\times 10^{-6}$ learning rate with 200 steps of linear warmup, followed by cosine annealing to the end. We use Axolotl  [^4] https://github.com/axolotl-ai-cloud/axolotl for our SFT experiments.
@@
As shown in Table 4, after adaptation, there is a huge performance drop mainly on generative tasks as the computation flow is messed up. But after light retraining, the hybrid Ladder Llama is able to reach the same level of performance with the original Llama. By applying Ladder Residual on the last 16 layers, we can obtain 21% end-to-end wall clock speed up for inference with TP world size of 8 and batch size of 1. Our results demonstrate the potential of Ladder Residual being a drop-in adaptation technique for an existing model to make it faster without sacrificing performance. We additionally experiment with applying Ladder Residual to the last 20 layers of Llama and found that it leads to a slight drop in performance. There is a chance that with longer adaptation, or smarter adaptation techniques like distillation or iterative training, we can obtain a Ladder-Llama with more layers adapted. We leave the further exploration to future work.
@@ 5 Related Work
Overlapping communication has been a widely explored area in prior works in order to achieve higher performance for distributed training. For Tensor Parallelism, prior works (Jangda et al., 2022; Wang et al., 2022; NVIDIA, 2023) decompose the communication into more fine-grained operations in order to find computations with no dependency to overlap. Our work doesn’t rely on such decompositions and therefore doesn’t require Sequence Parallelism to handle the partitioned activations before all-gather. In FSDP (FairScale authors, 2021), the all-gather communication is usually prefetched to be overlapped with the communication. Pipeline Parallelism (NVIDIA, 2023; Li et al., 2023a; Lamy-Poirier, 2023) on the other hand, chunks the data into mini-batches which creates more opportunity for overlapping. Compared to these other parallelism approaches, TP has the advantage to be independent of the batch size or sequence length, and can partition the computation as much as possible given enough GPUs in theory.
@@
Prior works have explored various alternative designs for Transformer, for example parallel attention and mlp (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021), linear attention (Katharopoulos et al., 2020), Grouped Query Attention (Ainslie et al., 2023), Cross-Layer Attention (Brandon et al., 2024) to improve the training and inference efficiency. Concurrently, Kraken (Prabhakar et al., 2024) treats each shard of weight as independent networks and only passes information between them occasionally to eliminate communication. However, the trained model under this design can only be used under a preset degree of parallelization. DistriFusion (Li et al., 2024) notices a similar phenomenon in the diffusion model, where the input is similar in adjacent diffusion steps, and overlaps the synchronization of distributed patches by reusing the feature maps from the previous diffusion step at the current step. Some of these variants are more widely adopted than others, due to the degree of impact they have on performance and efficiency. Past works have also considered adapting an existing checkpoint to these efficient variants.  Ainslie et al. (2023) extracted grouped-query attention from a multi-head attention model, and  Wang et al. (2024) converted a Llama model to a Mamba model by retraining on 50B tokens to close the performance gap. Comparatively, our adaptation is much lighter (3B tokens), showing that the representation shift introduced by Ladder Residual is easier to recover.  Wang et al. (2024) considered converting a Llama model to a Mamba model and used distillation to retrain the converted model. Such training paradigms that specifically tune the model to align with the original model could further improve the Ladder Residual based models.
@@ 6 Conclusion
We introduce Ladder Residual, architectural modifications that allow overlapping communication with computationfor model parallelism. We show that when running Tensor Parallelism, Ladder Residual can achieve great speed up across various model sizes, batch sizes, and number of GPUs. When applying Ladder Residual to Llama-3.1 8B Instruct, we only need lightweight retraining to reach the same level of performance as the original model while being 21% faster, showing its potential to be a plug-in for any pretrained Transformer. We also trained a 1B and 3B Ladder Transformer from scratch, and found that they are comparable to the standard Transformer of the same size while achieving over 55% speedup. Given that such a simple architectural change can obviate the need for expensive interconnects while maintaining model quality, we hope that our method will inspire even closer co-design between model architecture and inference systems.
7 Acknowledgement
We thank Tianle Cai, Devendra Chaplot, Igor Babuschkin, Albert Gu for helpful discussions and feedback that improved this work.